# OmnimatteZero: Training-free Real-time Omnimatte with Pre-trained Video Diffusion Models

### v1

@@ Abstract
Omnimatte aims to decompose a given video into semantically meaningful layers, including the background and individual objects along with their associated effects, such as shadows and reflections. Existing methods often require extensive training or costly self-supervised optimization. In this paper, we present OmnimatteZero, a training-free approach that leverages off-the-shelf pre-trained video diffusion models for omnimatte. It can remove objects from videos, extract individual object layers along with their effects, and composite those objects onto new videos. We accomplish this by adapting zero-shot image inpainting techniques for video object removal, a task they fail to handle effectively out-of-the-box. We then show that self-attention maps capture information about the object and its footprints and use them to inpaint the object’s effects, leaving a clean background. Additionally, through simple latent arithmetic, object layers can be isolated and recombined seamlessly with new video layers to produce new videos. Evaluations show that OmnimatteZero not only achieves superior performance in terms of background reconstruction but also sets a new record for the fastest Omnimatte approach, achieving real-time performance with minimal frame runtime.
1 Introduction
Image diffusion models have achieved remarkable success in image generation [26, 27, 24, 1, 3]. Beyond generation, these models have proven highly effective for various image editing tasks, including image inpainting [21, 6], object removal [32, 4], and image matting [11] with some methods achieving these edits without additional training  [21, 6].
With the recent introduction of video diffusion models [9, 14, 29] video editing tasks have become increasingly relevant. One particularly challenging variant of video editing is the video matting task, commonly referred to as Omnimatte [31, 19]. Omnimatte aims to decompose an input video into a background layer and multiple foreground layers, where each foreground layer contains an object of interest along with its associated effects, such as shadows and reflections. These extracted layers facilitate various video editing applications, including object removal, background replacement, retiming [18], and object duplication [15].
@@ 1 Introduction
Various video inpainting and Omnimatte methods have been proposed, but all rely on model training or optimization. Recent video inpainting approaches  [16, 36] are trained on annotated datasets for object removal but fail to account for their associated effects on the scene. Similarly, Omnimatte methods require expensive self-supervised optimization for each video [19, 28, 17, 15], with some demanding large-scale dataset training [32, 15] or 3D scene modeling using meshes or NeRF [28, 17]. These methods typically involve hours of training per video and prolonged inference times for rendering each frame.
@@
Given the success of fast, training-free approaches in image editing [21, 6], a natural question arises: can these methods be effectively adapted to video editing?
@@
In this paper, we present a new approach called OmnimatteZero, the first real-time training-free generative method to Omnimatte. OmnimatteZero (named for its zero training and optimization requirements) leverages off-the-shelf pre-trained video diffusion models to remove all objects from a video, extract individual object layers along with their associated effects, and finally composite them onto new backgrounds. This entire process is performed efficiently during diffusion inference, making it significantly faster than current methods. See Figure LABEL:fig:fig1 for illustration.
Our method begins by adapting a training-free image inpainting approaches [21, 22] for object inpainting using video diffusion models. We demonstrate that directly applying these methods to videos does not work and propose a simple yet effective solution for object removal in videos. Next, we show that self-attention maps capture information about both the object and its traces in the scene, such as shadows and reflections. We use these maps to inpaint the object’s effects, resulting in a clean background. For foreground layer extraction, we find that the object layer, along with its footprints, can be simply extracted by linearly subtracting the latent of the video (with the object) from the clean background latent (without the object). This object layer can then be recomposed to a new background video, by adding the latent encodings, enabling fast and easy layer insertion and composition.
@@
To validate our approach, we apply it to two video diffusion models, LTXVideo [9] and Wan2.1 [29], and evaluate it on standard Omnimatte benchmarks both quantitatively and qualitatively. Our method outperforms all supervised and self-supervised existing inpainting and Omnimatte techniques across all benchmarks in terms of background reconstruction and object layer extraction. Additionally, since LTXVideo generates videos in real-time, we set a new record for the fastest Omnimatte approach, achieving real-time performance with minimal runtime per frame.
2 Related Work
2.1 Omnimatte
@@
The first Omnimatte approaches [12, 19] use motion cues to decompose a video into RGBA matte layers, assuming a static background with planar homographies. Optimization isolates dynamic foreground layers, taking 3 to 8.5 hours per video and $\sim$2.5 seconds per frame to render. Later methods enhanced Omnimatte with deep image priors [20, 8] and 3D representations [17, 28, 33], but still rely on strict motion assumptions and require extensive training ($\sim$6 hours) with rendering times of 2.5-3.5 seconds per frame. A recent generative approach [15] uses video diffusion priors to remove objects and effects, with test-time optimization for frame reconstruction, taking $\sim$9 seconds per frame. In this paper, we propose a novel, training and optimization-free method using off-the-shelf video diffusion models, achieving real-time performance in 0.04 seconds per frame.
2.2 Video Inpainting
@@
Earlier video inpainting methods [5, 30, 10] used 3D CNNs for spatiotemporal modeling but struggled with long-range propagation due to limited receptive fields. Pixel-flow approaches [7, 35, 34] improved texture and detail restoration by utilizing adjacent frame information. Recently, Zhou et al. [36] introduced ProPainter, which enhances reconstruction accuracy and efficiency by leveraging optical flow-based propagation and attention maps. Li et al. [16] proposed DiffuEraser, which combines a pre-trained text-to-image diffusion model with BrushNet, a feature extraction module, to generate temporally consistent videos. They use a two-stage process where spatial inpainting is extended to the spatiotemporal domain in a separate step. In this paper, we propose a single-step object inpainting approach, directly applied during video diffusion model inference, implicitly leveraging spatial and temporal information from adjacent frames for inpainting.
2.3 Diffusion Transformers for Video Generation
@@
Recent advancements in diffusion models have significantly improved text-to-video generation. These models, trained on large datasets of video-text pairs, can generate visually compelling videos from text descriptions. Recent work introduced LTX-Video [9], a real-time transformer-based latent diffusion model capable of generating videos at over 24 frames per second. It builds on 3D Video-VAE with a spatiotemporal latent space, a concept also used in Wan2.1 [29] and the larger HunyuanVideo [14]. Unlike traditional methods that treat the video VAE and denoising transformer separately, LTX-Video integrates them within a highly compressed latent space (1:192 compression ratio), optimizing their interaction. This spatiotemporal downscaling results in $32\times 32\times 8$ pixels per token while maintaining high-quality video generation. Similarly, HunyuanVideo employs a large-scale video generation framework, integrating an optimized VAE and diffusion-based transformer to achieve excellent visual quality and motion consistency. These models demonstrate the trend of combining high-compression VAEs with transformer-based diffusion models for efficient and scalable video generation.
Figure 1: Overview of OmnimatteZero-From left to right: The video and its corresponding object-masked version are encoded into the latent space. The spatio-temporal footprint of the target object is then identified and extracted from this encoding. Self-attention maps are leveraged to recognize effects associated with the object (such as a cat’s reflection), which are incorporated into both the object mask and the latent encoding. Using this mask, we apply two imperfect video inpainting methods: (a) Background-Preserving Inpainting, which retains background details but may leave behind traces of the object, and (b) Object-Removing Inpainting, which eliminates the object but might introduce background distortions. We refine the results through Attention-based Latent Blending, selectively combining the strengths of each inpainting method. Finally, few denoising steps of the video diffusion model enhances the blended output, producing a high-quality reconstruction with the object removed and the background well-preserved, as indicated by high PSNR values.
3 Motivation: The failure of image inpainting approaches for videos
We begin by exploring training-free approaches to image diffusion inpainting and extend it directly to video diffusion models. Our starting point is RePaint [21], a widely used and straightforward inpainting method. RePaint iteratively reconstructs masked regions by adding Gaussian noise to the background, aligning it with the denoising process. A denoising step then generates missing content while conditioning on known areas. This cycle repeats, refining the output by reapplying noise and rerunning the process. When directly applying RePaint to videos, our goal is to use it for inpainting both spatially and temporally (across the entire latent space, rather than frame by frame), as frame-by-frame inpainting would lead to temporal inconsistencies between frames. However, we observe significant limitations with this approach. While the background remains well-aligned with the input video, the inpainted regions often degrade into incoherent blobs, lacking both structural fidelity and temporal consistency across frames (see Figure 1 Visualization (a), red dashed blob). We hypothesize that this failure stems from the fundamental differences between image and video inpainting. In image inpainting, the model only needs to reconstruct missing regions using spatial information to ensure coherence within a the image. In contrast, video inpainting requires consistency both spatially and temporally, maintaining smooth transitions across consecutive frames. While in-frame (single image) inpainting strongly relies on hallucination, in video, missing information can be inferred from different spatio-temporal regions. These additional constraints make video inpainting much more challenging, as the model must not only generate plausible spatial content but also ensure accuracy and maintain temporal stability both within and outside the masked region, an aspect that RePaint fails to achieve.
To validate our hypothesis, we explore another widely used zero-shot inpainting approach known as Img2Img or SDEdit [22]. This method adds noise to the image latent, applying higher noise to the masked region and lower noise to the background to control changes. The model then iteratively denoises the noisy latent, aiming to modify the masked region. This method does not guarantee perfect background preservation and often introduces unintended artifacts into the background.
When extending this approach to video diffusion models and applying spatio-temporal denoising, the model effectively removes objects and plausibly inpaints the masked regions. However, this process further generates noise and artifacts into the background (see Figure 1, Visualization (b)), causing the degradation of the outcome. This observation confirms that video diffusion models can remove objects from video content, but only when the requirement for high fidelity background preservation is relaxed. However, this comes at the cost of introducing artifacts and noise in the resulting video.
4 Method: OmnimatteZero
We start with the process of object removal and inpainting. Building on our previous observations, we strive to develop a method that ensures the background remains consistent with the original video while also accurately inpainting the missing object. Furthermore, we seek to eliminate the effects the object has on the scene (e.g. shadows, reflections), a challenge where existing inpainting methods fall short.
Our model takes as input a video along with an object mask for each frame, following the common Omnimatte setup [19]. Alternatively, a single-frame mask can be used and propagated across all frames using a general-purpose video segmentation model, such as SAM2 [25]. We denote this mask as $M^{p}_{obj}\in\{0,1\}^{F\times W\times H}$, with $W,H$ denoting the frame size and $F$ the number of frames, and $p$ designates pixel space.
To address the challenges faced by current zero-shot inpainting method, we propose a simple yet effective approach. The method is illustrated in Figure 1. We start with an input video with RGB frames, $V\in{\mathbb{R}}^{F\times W\times H\times 3}$, The video is then passed through the VAE yielding a compressed latent feature map $Z\in{\mathbb{R}}^{f\times w\times h\times c}$, with $w,h,c$ denoting the latent space dimensions and $f$ represents the number of temporally compressed latent frames.
Applying a mask directly in latent space is challenging due to the unclear mapping from pixel space to the latent space. We start by taking the object masks (per-frame) and computing a binary video, where the frames are binary images, and pass this video as an RGB input (with duplicated channels) through the VAE encoder as well. This process produces a latent representation where high values indicate the masked object, enabling its spatio-temporal footprint identification via simple thresholding. We denote this mask as $M_{obj}\in\{0,1\}^{f\times w\times h}$ (see Figure 1 at left).
Next, we apply two inpainting processes, simultaneously: (1) Background-Preserving video inpainting, which attempt to remove the object while maintaining the background (inspired by [21]), $\mathcal{F}^{BP}:V\rightarrow Z^{BP}$, and (2) Object-Removal inpainting, which successfully removes the object but fails to preserve the background (inspired by [22]), $\mathcal{F}^{OR}:V\rightarrow Z^{OR}$. We apply both inpaintings for $k$ denoising steps. At step $k$, we blend them as:
$Z_{k}=M_{obj}\odot Z^{BP}_{k}+(1-M_{obj})\odot Z^{OR}_{k}$ (1)
where $\odot$ denotes element-wise multiplication. Next, we perform an additional $T_{steps}-k$ denoising steps with the video diffusion model, completing the remaining diffusion steps, to blend the two latent representations. This ensures a smooth transition between the inpainted object region and the surrounding background and results in a video where the object is removed while the background closely matches the original input (Figure 1 Visualization (c)).
Up to this point, we have demonstrated how objects can be removed from videos. However, this process does not eliminate their associated effects. In the following chapters, we show how to identify and remove both the object and its effects, extract object layers, and composite them with new layers to generate new videos.
4.1 Removing associated object effects

### v2

# OmnimatteZero: Fast Training-free Omnimatte with Pre-trained Video Diffusion Models

@@ Abstract.
In Omnimatte, one aims to decompose a given video into semantically meaningful layers, including the background and individual objects along with their associated effects, such as shadows and reflections. Existing methods often require extensive training or costly self-supervised optimization. In this paper, we present OmnimatteZero, a training-free approach that leverages off-the-shelf pre-trained video diffusion models for omnimatte. It can remove objects from videos, extract individual object layers along with their effects, and composite those objects onto new videos. These are accomplished by adapting zero-shot image inpainting techniques for video object removal, a task they fail to handle effectively out-of-the-box. To overcome this, we introduce temporal and spatial attention guidance modules that steer the diffusion process for accurate object removal and temporally consistent background reconstruction. We further show that self-attention maps capture information about the object and its footprints and use them to inpaint the object’s effects, leaving a clean background. Additionally, through simple latent arithmetic, object layers can be isolated and recombined seamlessly with new video layers to produce new videos. Evaluations show that OmnimatteZero not only achieves superior performance in terms of background reconstruction but also sets a new record for the fastest Omnimatte approach, achieving real-time performance with minimal frame runtime. Project Page.
Omnimatte, Video Diffusion Model, Training-free, Real-time
journal: TOGjournalvolume: 00journalnumber: 0article: 000publicationmonth: 8submissionid: papers_1753
1. Introduction
Extracting or isolating objects in videos is a central problem in video understanding and editing, with important applications such as removing, modifying, or enhancing elements within a video scene. One important flavor of this task is video matting (Chuang et al., 2002), which involves decomposing a video into a background layer and one or more foreground layers, each representing an individual object. A particularly challenging form of video matting is Omnimatte (Wang and Adelson, 1994; Lu et al., 2021), where each foreground object is not only isolated but also includes associated effects such as shadows, reflections, and scene perturbations. The layers extracted by video matting facilitate various video editing applications, including object removal, background replacement, retiming (Lu et al., 2020), and object duplication (Lee et al., 2024), many of which require estimating occluded background regions using techniques like video inpainting or omnimatting.
@@ 1. Introduction
However, despite their promise, existing methods for video inpainting and Omnimatte face significant limitations. Recent inpainting approaches (Li et al., 2025; Zhou et al., 2023), while trained on annotated datasets for object removal, fail to account for the nuanced effects that objects impart on the scene. Similarly, current Omnimatte methods demand computationally intensive self-supervised optimization for each video (Lu et al., 2021; Suhail et al., 2023; Lin et al., 2023; Lee et al., 2024). Some approaches further rely on large hand-crafted training datasets (Winter et al., 2024; Lee et al., 2024) or require 3D scene modeling using meshes or NeRFs (Suhail et al., 2023; Lin et al., 2023), resulting in hours of training and slow per-frame rendering.
@@
Given the success of fast, training-free approaches in image editing (Lugmayr et al., 2022; Corneanu et al., 2024), a natural question arises: can these methods be effectively adapted to video editing?
@@
In this paper, we present a new approach called OmnimatteZero, the first training-free generative method to Omnimatte. OmnimatteZero (named for its zero training and optimization requirements) leverages off-the-shelf pre-trained video diffusion models to remove all objects from a video, extract individual object layers along with their associated effects, and finally composite them onto new backgrounds. This entire process is performed efficiently during diffusion inference, making it significantly faster than current methods. See Figure 1 for illustration.
We begin by adapting training-free image inpainting methods (Lugmayr et al., 2022; Meng et al., 2022) to handle object inpainting in videos using diffusion models. However, we find that simply applying these frame-by-frame techniques to videos leads to poor results: missing regions are not filled convincingly, and temporal consistency breaks down. To address this, we propose two attention guidance modules that can be easily integrated into pre-trained video diffusion models. These modules guide the diffusion process to achieve both accurate object removal and consistent background reconstruction. The first module, Temporal Attention Guidance (TAG), uses background patches from nearby frames. By modifying self-attention layers, it encourages the model to use these patches to reconstruct masked regions, improving temporal coherence. The second module, Spatial Attention Guidance (SAG), operates within each frame to refine local details and enhance visual plausibility.
We also observe that self-attention maps contain cues not just about the object but also its indirect effects, such as shadows and reflections. By leveraging this information, our method can inpaint these subtle traces, yielding cleaner backgrounds. For foreground layer extraction, we find that the object (and its traces) can be isolated by subtracting the latent representation of the clean background from the original video latent. This extracted object layer can then be seamlessly inserted into new backgrounds by simply adding the latents, enabling fast and flexible video composition.
@@
To validate our approach, we apply it to two video diffusion models, LTXVideo (HaCohen et al., 2024) and Wan2.1 (video, 2025), and evaluate it on standard Omnimatte benchmarks both quantitatively and qualitatively. Our method outperforms all supervised and self-supervised existing inpainting and Omnimatte techniques across all benchmarks in terms of background reconstruction and object layer extraction. Notably, OmnimatteZero requires no training or additional optimization, functioning entirely within the denoising process of diffusion models. This positions OmnimatteZero as the fastest Omnimatte technique to date.
2. Related Work
2.1. Omnimatte
@@
The first Omnimatte approaches (Kasten et al., 2021; Lu et al., 2021) use motion cues to decompose a video into RGBA matte layers, assuming a static background with planar homographies. Optimization isolates dynamic foreground layers, taking 3 to 8.5 hours per video and $\sim$2.5 seconds per frame to render. Later methods enhanced Omnimatte with deep image priors (Lu et al., 2022; Gu et al., 2023) and 3D representations (Lin et al., 2023; Suhail et al., 2023; Wu et al., 2024), but still rely on strict motion assumptions and require extensive training ($\sim$6 hours) with rendering times of 2.5-3.5 seconds per frame. A recent generative approach (Lee et al., 2024) uses video diffusion priors to remove objects and effects, with test-time optimization for frame reconstruction, taking $\sim$9 seconds per frame. In this paper, we propose a novel, training and optimization-free method using off-the-shelf video diffusion models, achieving real-time performance in 0.04 seconds per frame.
2.2. Video Inpainting
@@
Earlier video inpainting methods (Chang et al., 2019; Wang et al., 2019; Hu et al., 2020) used 3D CNNs for spatiotemporal modeling but struggled with long-range propagation due to limited receptive fields. Pixel-flow approaches (Gao et al., 2020; Zhang et al., 2022a, b) improved texture and detail restoration by utilizing adjacent frame information. Recently, Zhou et al. (2023) introduced ProPainter, which enhances reconstruction accuracy and efficiency by leveraging optical flow-based propagation and attention maps. VideoPainter (Bian et al., 2025) proposed a dual-stream video inpainting method that injects video context into a pre-trained video diffusion model using a lightweight encoder. (Li et al., 2025) proposed DiffuEraser, which combines a pre-trained text-to-image diffusion model with BrushNet, a feature extraction module, to generate temporally consistent videos. In this paper, we propose a single-step object inpainting approach, directly applied during video diffusion model inference, implicitly leveraging spatial and temporal information from adjacent frames for inpainting.
2.3. Diffusion Transformers for Video Generation
@@
Recent advancements in diffusion models have significantly improved text-to-video generation. These models, trained on large datasets of video-text pairs, can generate visually compelling videos from text descriptions. Recent work introduced LTX-Video (HaCohen et al., 2024), a real-time transformer-based latent diffusion model capable of generating videos at over 24 frames per second. It builds on 3D Video-VAE with a spatiotemporal latent space, a concept also used in Wan2.1 (video, 2025) and the larger HunyuanVideo (Kong, 2024). Unlike traditional methods that treat the video VAE and denoising transformer separately, LTX-Video integrates them within a highly compressed latent space (1:192 compression ratio), optimizing their interaction. This spatiotemporal downscaling results in $32\times 32\times 8$ pixels per token while maintaining high-quality video generation. Similarly, HunyuanVideo employs a large-scale video generation framework, integrating an optimized VAE and diffusion-based transformer to achieve excellent visual quality and motion consistency. These models demonstrate the trend of combining high-compression VAEs with transformer-based diffusion models for efficient and scalable video generation.
[…]
[motivation figure]motivation figure
Figure 2. Comparison of object removal results using (a) vanilla image inpainting extended to video, (b) Vid2Vid zero-shot inpainting, and (c) our guidance-based approach. Vanilla methods fail to maintain temporal consistency or clean background reconstruction, while our method achieves coherent, artifact-free inpainting across frames.
[…]
[method figure]method figure
Figure 3. Overview of our Object Removal strategy in OmnimatteZero. (a) We first identify potential background correspondences across frames. (b) Temporal Attention Guidance (TAG): Temporal attention scores between a foreground point and its background correspondences are replaced with the average attention between all background pairs, promoting consistent inpainting across time. (c) Spatial Attention Guidance (SAG): Within a frame, the attention from a foreground point to nearby background points is adjusted to reflect the mean attention among background points themselves, improving inpainting quality when temporal context is unavailable.
3. Motivation: The failure of image inpainting approaches for videos
We begin by extending training-free image-diffusion inpainting techniques to video diffusion models. Our reference point is vanilla image inpainting (Lugmayr et al., 2022), a straightforward method that starts with a random noise at the masked area and, at each diffusion step $t$, injects Gaussian noise $\epsilon_{t}$ into the original background area so the denoiser can reconstruct the masked region while preserving the background during denoising. For video, we apply it across the full video latent to maintain both spatial and temporal coherence, since frame-by-frame inpainting leads to temporal artifacts like flickering. In practice, this vanilla approach does preserve the original background alignment, however causes the masked areas to degrade into incoherent, temporally unstable blobs (Fig. 2a). We attribute this breakdown to a core difference between image and video models: image inpainting ensures spatial consistency within a single frame, while video inpainting must also maintain temporal consistency across frames. Unlike image inpainting, which relies heavily on hallucination, video inpainting must infer missing content from both spatial and temporal context, making it significantly more challenging. We further examine a second zero-shot inpainting method: Vid2Vid (HaCohen et al., 2024) (an adaptation from Img2Img (Meng et al., 2022)). This method applies heavy noise inside the mask and lighter noise elsewhere in the latent video, then denoises the entire latent. Vid2Vid effectively removes the target object and synthesizes plausible fills, however, it also introduces artifacts into the background and motion (Fig.2b).
While recent diffusion-based solutions (Bian et al., 2025; Lee et al., 2024) overcome these issues by training dedicated video diffusion models for object removal, we pose the following question: Can we develop a training free method that simultaneously (i) convincingly fills the masked region, (ii) preserves the background so that it is indistinguishable from the original video, and (iii) eliminates all residual evidence of the object’s presence, including secondary effects like shadows and reflections?.
4. Preliminaries: Spatio–Temporal Attention in Video Diffusion Transformers
Video diffusion transformers interleave spatio-temporal self-attention modules in each diffusion block to model both fine-grained structure within frames and consistent dynamics across time. Given a video, it’s latent tensor is $Z\in\mathbb{R}^{f\times w\times h\times c}$, where $f$ is the number of temporally compressed frames, and $w$, $h$, and $c$ are the width, height, and channel dimensions in latent space. We reshape $Z$ into $Z\in\mathbb{R}^{(f\times n)\times c}$, where $n=h\times w$ is the number of tokens per frame. Self‐attention is then applied over all $f\times n$ tokens:
$A\;=\;\mathrm{Attention}\bigl{(}Q(z),\,K(z)\bigr{)}\;\in\;\mathbb{R}^{(f\times n)\times(f\times n)}.$
where $Q,K\in\mathbb{R}^{w\times h\times c}$ are the queries and keys, respectively. This matrix $A$ encodes both spatial and temporal interactions in one operation. We denote tokens by $(i,p)$ with $i\!=\!1,\dots,f$ the frame index and $p\!=\!1,\dots,n$ the spatial position. Spatial attention within frame $i$ (i.e. interactions among tokens $(i,p)$ and $(i,q)$) is given by the in-frame weights $A_{\mathrm{spatial}}^{(i)}(p,q)\;=\;A_{(i,p),(i,q)}.$ Temporal attention between frame $i$ to frame $j\neq i$ (i.e. interactions between tokens $(i,p)$ and $(j,q)$) is given by $A_{\mathrm{temporal}}^{(i,j)}(p,q)\;=\;A_{(i,p),(j,q)}.$
5. Method: OmnimatteZero
In this section, we first describe how given a video $\mathcal{V}\in\mathbb{R}^{F\times W\times H\times 3}$ and an object mask for each frame $M_{obj}\in\{0,1\}^{F\times H\times W}$, one can use off-the-shelf video diffusion models to remove the target object (Section 5.1) and its associated visual effects (Section 5.2). Next, we detail our strategy for isolating foreground objects together with their residual effects (Section 5.3). Finally, we demonstrate how to recombine these object–effect layers onto arbitrary background videos (Section 5.4), enabling flexible and realistic video editing.
5.1. Object Removal
Zero-shot video inpainting with diffusion models  (Meng et al., 2022; Lugmayr et al., 2022) frequently struggles because the model must hallucinate content that is coherent both spatially and temporally ( Section 3). Recent studies show that the self-attention mechanism within diffusion models is crucial for maintaining coherence during model generation (Luo et al., 2025). However, it remains unclear how self-attention can effectively guide the inpainting process.
Our core observation is that in real-world videos, the necessary background information is often readily available in neighboring frames. This principle has been leveraged in traditional video editing techniques (Wexler et al., 2004; Criminisi et al., 2004). However, they often produce non-realistic results or do not scale for long videos. Motivated by this, we introduce a method that explicitly guides a video diffusion model to reconstruct missing pixels by leveraging spatial and temporal context, rather than relying on the model to infer this implicitly. To further enhance spatial inpainting within individual frames, particularly when background details are absent from adjacent frames (e.g., due to static scenes or lack of camera/object motion), we exploit contextual cues from regions surrounding the masked area within the same frame. These temporal and spatial guidance are facilitated through careful manipulations of spatial and temporal self-attention.
We thus introduce two novel modules: Temporal Attention guidance (TAG) and Spatial Attention Guidance (SAG). These modules seamlessly integrate into existing pre-trained video diffusion models, effectively steering the diffusion process toward object removal and coherent background reconstruction.
Identifying Potential Background Patches from Neighboring Frames. To accurately inpaint masked areas in a specific frame, it is essential to identify corresponding background patches from neighboring frames, while preserving awareness of the video’s underlying 3D structure. This task can be effectively addressed using Track-Any-Point (TAP-Net) (Doersch et al., 2023), a real-time tracking framework. TAP-Net captures spatial and temporal coherence by tracking points from one frame to their corresponding locations across subsequent frames, as illustrated in Figure 3a, where points from frame $f_{i}$ are matched across other frames. Formally, given a point $p$ on the object in frame $i$, TAP-Net identifies its corresponding point set among frames $C=\{C_{j}|j\neq i\}$. To ensure that only background regions are considered, we discard points in $C$ that intersect the object mask $\mathcal{M}$ in their respective frames.
Temporal Attention Guidance (TAG): To effectively guide the denoising process toward inpainting using neighboring frames, we explicitly set the temporal attention score between a foreground point $p$ (in frame $i$) and each of its background correspondences $c\in C$ to be the mean temporal attention observed between all distinct pairs of corresponding background points (see Figure 3b). Formally, we first compute the mean temporal attention
(1) $\bar{A}_{temporal}=\frac{1}{|C|(|C|-1)}{\sum_{j=1}^{|C|}{\sum_{\begin{subarray}{c}k=1\\ k\neq j\end{subarray}}^{|C|}{A^{(j,k)}_{temporal}(c_{j},c_{k})}}}.$
Since the attention is bidirectional, the denominator $|C|(|C|-1)$ correctly enumerates all ordered pairs $(c_{j},c_{k})$ where $j\neq k$, ensuring that each directional interaction is counted exactly once.
We then replace the temporal-attention score from $p$ to every background point $c$ as follows:
$A_{temporal}^{(i,j)}(p,c_{j})=\bar{A}_{temporal},\quad\mbox{\Large$\forall$}c_{j}\in C.$
This assignment transfers a consensus cue from the entire background set to the foreground point, encouraging consistent inpainting across frames.
Spatial Attention Guidance (SAG): To further improve spatial inpainting within each frame and to address cases where a point lacks correspondences in other frames, we reinforce spatial attention within the same frame. Specifically, for any foreground point $p$ in frame $i$, we set its spatial attention scores with surrounding background points $S$ to be the mean spatial attention computed among all pairs of background points within the same frame (see Figure 3c). Formally:
(2) $\bar{A}_{spatial}=\frac{1}{|S|(|S|-1)}{\sum_{j=1}^{|S|}{\sum_{\begin{subarray}{c}k=1\\ k\neq j\end{subarray}}^{|S|}{A^{(I)}_{spatial}(s_{j},s_{k})}}}.$
Finally, we replace the spatial-attention score from $p$ to every background point $s$ as follows:
$A^{i}_{\text{spatial}}(p,s)=\bar{A}_{spatial},\quad\mbox{\Large$\forall$}s\in S.$
This strategy ensures coherent spatial inpainting even in the absence of temporal information.
5.2. Removing associated object effects